OpenStudy (anonymous):

Let f(x) = 2x+3 and g(x) = x2 Part 1 Find f(g(x)) Part 2 Find g(f(x)) Part 3 Use complete sentences to explain any difference in part 1 and part 2.

Parth (parthkohli):

Just substitute the functions with the values, for example: \(\Large \color{purple}{\rightarrow f(g(x)) = f(x^2) = 2(x^2) + 3 }\)